1. The incidence of Miller syndrome, a rare, recessive Mendelian disease, is 0.32%. Assuming that the population is in Hardy-Weinberg equilibrium for the gene that causes this disease, calculate the percent of the population that is a carrier for Miller syndrome.

2. Due to assortive mating, the frequency of genotypes is AA: 46%; Aa: 0%; aa: 54%

a.Calculate p and q for this population.

b.What is the expected genotype frequencies in Hardy-Weinberg equilibriumc. Assume that assortive mating ceases. How many generations will it take for the genotype frequencies to relax to Hardy-Weinberg equilibrium? Hint: calculate the expectant genotype frequencies in the next generation by identifying the possible mating classes (e.g. AA x AA), their probabilities, and their possible offspring. Sum these up and compare to Hardy-Weinberg equilibrium.
Answer #1

1.

Normal = MM, Mm

Affected = mm

mm = 0.32%

m = √0.32 = 0.5656

M = 1 - m = 1 - 0.5656 = 0.4344

Frequency of heterozygotes = Mm = 2 × 0.5656 × 0.4344 = 0.4913

Percentage of heterozygotes (carriers) = 0.4913 × 100 = 49.13%

Please rate high.